OpenStudy (lyssanicoleee):

In a certain mid-size city, 90% of the high school graduates are from public schools and the other 10% are from private schools. Of the private school grads, 80% go on to complete college, but only 50% of the public school grads complete college. A. What is the probability that a high school graduate from this city will complete college?

OpenStudy (lyssanicoleee):

B. What is the probability that a person from this city who completes college is a graduate of a private high school?

OpenStudy (kropot72):

Let A be the event 'public school graduate completes college'. P(A) = 0.9 * 0.5 Let B be the event 'private school graduate completes college'. P(A) = 0.1 * 0.8 Let C be the event 'high school graduate from that city will complete college'. P(C) = P(A) + P(B) @lyssanicoleee Can you now finish this?